### v1

@@ 6 CONCLUSION
Using DesignQA, we conducted a rigorous evaluation on variants of two state-of-the-art MLLMs: GPT4 and LLaVA. We uncovered significant limitations in their current abilities to accurately interpret complex technical documents, specifically in referencing relevant requirements and extracting numerical information from technical images. Our research highlights the need for advancements in AI models to enhance their comprehension of engineering requirements and documentation, suggesting directions for future efforts. Our work aims to bridge the gap in AI’s capability to support engineering design processes more effectively, paving the way for sophisticated AI-assisted engineering solutions.
@@ Acknowledgments
The authors wish to express their gratitude to Henry Smith and the MIT Motorsports team for their critical contribution to dataset creation and verification, which significantly enhanced the quality of this research. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
Список литературы
  • [1] OpenAI. ‘‘GPT-4V(ision) System Card.’’ 2023. URL https://api.semanticscholar.org/CorpusID:263218031.
  • [2] Vaswani, Ashish, Shazeer, Noam, Parmar, Niki, Uszkoreit, Jakob, Jones, Llion, Gomez, Aidan N, Kaiser, Łukasz and Polosukhin, Illia. ‘‘Attention is all you need.’’ Advances in neural information processing systems Vol. 30 (2017).
  • [3] Bubeck, Sébastien, Chandrasekaran, Varun, Eldan, Ronen, Gehrke, Johannes, Horvitz, Eric, Kamar, Ece, Lee, Peter, Lee, Yin Tat, Li, Yuanzhi, Lundberg, Scott et al. ‘‘Sparks of artificial general intelligence: Early experiments with GPT-4. arXiv.’’ arXiv preprint arXiv:2303.12712 (2023).
  • [4] Barbhuiya, Rejaul Karim. ‘‘Introduction to Artificial Intelligence: Current Developments, Concerns and Possibilities for Education.’’ Indian Journal of Educational Technology Vol. 5 No. 2 (2023): p. 266.
  • [5] Thirunavukarasu, Arun James, Ting, Darren Shu Jeng, Elangovan, Kabilan, Gutierrez, Laura, Tan, Ting Fang and Ting, Daniel Shu Wei. ‘‘Large language models in medicine.’’ Nature medicine Vol. 29 No. 8 (2023): pp. 1930–1940.
  • [6] Clusmann, Jan, Kolbinger, Fiona R, Muti, Hannah Sophie, Carrero, Zunamys I, Eckardt, Jan-Niklas, Laleh, Narmin Ghaffari, Löffler, Chiara Maria Lavinia, Schwarzkopf, Sophie-Caroline, Unger, Michaela, Veldhuizen, Gregory P et al. ‘‘The future landscape of large language models in medicine.’’ Communications Medicine Vol. 3 No. 1 (2023): p. 141.
  • [7] Kasneci, Enkelejda, Seßler, Kathrin, Küchemann, Stefan, Bannert, Maria, Dementieva, Daryna, Fischer, Frank, Gasser, Urs, Groh, Georg, Günnemann, Stephan, Hüllermeier, Eyke et al. ‘‘ChatGPT for good? On opportunities and challenges of large language models for education.’’ Learning and individual differences Vol. 103 (2023): p. 102274.
  • [8] Makatura, Liane, Foshey, Michael, Wang, Bohan, HähnLein, Felix, Ma, Pingchuan, Deng, Bolei, Tjandrasuwita, Megan, Spielberg, Andrew, Owens, Crystal Elaine, Chen, Peter Yichen et al. ‘‘How Can Large Language Models Help Humans in Design and Manufacturing?’’ arXiv preprint arXiv:2307.14377 (2023).
  • [9] Picard, Cyril, Edwards, Kristen M, Doris, Anna C, Man, Brandon, Giannone, Giorgio, Alam, Md Ferdous and Ahmed, Faez. ‘‘From Concept to Manufacturing: Evaluating Vision-Language Models for Engineering Design.’’ arXiv preprint arXiv:2311.12668 (2023).
  • [10] Ulrich, Karl T and Eppinger, Steven D. Product design and development. McGraw-hill (2016).
  • [11] Zeng, Yan, Zhang, Hanbo, Zheng, Jiani, Xia, Jiangnan, Wei, Guoqiang, Wei, Yang, Zhang, Yuchen and Kong, Tao. ‘‘What Matters in Training a GPT4-Style Language Model with Multimodal Inputs?’’ arXiv preprint arXiv:2307.02469 (2023).
  • [12] Wang, Lei, Hu, Yi, He, Jiabang, Xu, Xing, Liu, Ning, Liu, Hui and Shen, Heng Tao. ‘‘T-SciQ: Teaching Multimodal Chain-of-Thought Reasoning via Large Language Model Signals for Science Question Answering.’’ arXiv preprint arXiv:2305.03453 (2023).
  • [13] Liu, Haotian, Li, Chunyuan, Wu, Qingyang and Lee, Yong Jae. ‘‘Visual instruction tuning.’’ Advances in neural information processing systems Vol. 36 (2024).
  • [14] Shaham, Uri, Ivgi, Maor, Efrat, Avia, Berant, Jonathan and Levy, Omer. ‘‘ZeroSCROLLS: A Zero-Shot Benchmark for Long Text Understanding.’’ arXiv preprint arXiv:2305.14196 (2023).
  • [15] Xiong, Wenhan, Liu, Jingyu, Molybog, Igor, Zhang, Hejia, Bhargava, Prajjwal, Hou, Rui, Martin, Louis, Rungta, Rashi, Sankararaman, Karthik Abinav, Oguz, Barlas et al. ‘‘Effective long-context scaling of foundation models.’’ arXiv preprint arXiv:2309.16039 (2023).
  • [16] Richardson, Matthew, Burges, Christopher JC and Renshaw, Erin. ‘‘Mctest: A challenge dataset for the open-domain machine comprehension of text.’’ Proceedings of the 2013 conference on empirical methods in natural language processing: pp. 193–203. 2013.
  • [17] Rajpurkar, Pranav, Zhang, Jian, Lopyrev, Konstantin and Liang, Percy. ‘‘Squad: 100,000+ questions for machine comprehension of text.’’ arXiv preprint arXiv:1606.05250 (2016).
  • [18] Yang, Yi, Yih, Wen-tau and Meek, Christopher. ‘‘Wikiqa: A challenge dataset for open-domain question answering.’’ Proceedings of the 2015 conference on empirical methods in natural language processing: pp. 2013–2018. 2015.
  • [19] Dasigi, Pradeep, Lo, Kyle, Beltagy, Iz, Cohan, Arman, Smith, Noah A and Gardner, Matt. ‘‘A dataset of information-seeking questions and answers anchored in research papers.’’ arXiv preprint arXiv:2105.03011 (2021).
  • [20] Fu, Chaoyou, Chen, Peixian, Shen, Yunhang, Qin, Yulei, Zhang, Mengdan, Lin, Xu, Qiu, Zhenyu, Lin, Wei, Yang, Jinrui, Zheng, Xiawu et al. ‘‘MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models.’’ arXiv preprint arXiv:2306.13394 (2023).
  • [21] Liu, Yuan, Duan, Haodong, Zhang, Yuanhan, Li, Bo, Zhang, Songyang, Zhao, Wangbo, Yuan, Yike, Wang, Jiaqi, He, Conghui, Liu, Ziwei et al. ‘‘Mmbench: Is your multi-modal model an all-around player?’’ arXiv preprint arXiv:2307.06281 (2023).
  • [22] Lu, Pan, Mishra, Swaroop, Xia, Tanglin, Qiu, Liang, Chang, Kai-Wei, Zhu, Song-Chun, Tafjord, Oyvind, Clark, Peter and Kalyan, Ashwin. ‘‘Learn to explain: Multimodal reasoning via thought chains for science question answering.’’ Advances in Neural Information Processing Systems Vol. 35 (2022): pp. 2507–2521.
  • [23] Chen, Yang, Hu, Hexiang, Luan, Yi, Sun, Haitian, Changpinyo, Soravit, Ritter, Alan and Chang, Ming-Wei. ‘‘Can Pre-trained Vision and Language Models Answer Visual Information-Seeking Questions?’’ arXiv preprint arXiv:2302.11713 (2023).
  • [24] Song, Binyang, Zhou, Rui and Ahmed, Faez. ‘‘Multi-modal machine learning in engineering design: A review and future directions.’’ Journal of Computing and Information Science in Engineering Vol. 24 No. 1 (2024): p. 010801.
  • [25] Dima, Alden, Lukens, Sarah, Hodkiewicz, Melinda, Sexton, Thurston and Brundage, Michael P. ‘‘Adapting natural language processing for technical text.’’ Applied AI Letters Vol. 2 No. 3 (2021): p. e33.
  • [26] Brundage, Michael P, Sexton, Thurston, Hodkiewicz, Melinda, Dima, Alden and Lukens, Sarah. ‘‘Technical language processing: Unlocking maintenance knowledge.’’ Manufacturing Letters Vol. 27 (2021): pp. 42–46.
  • [27] Jiang, Shuo, Hu, Jie, Magee, Christopher L and Luo, Jianxi. ‘‘Deep learning for technical document classification.’’ IEEE Transactions on Engineering Management (2022).
  • [28] Zhu, Qihao and Luo, Jianxi. ‘‘Generative transformers for design concept generation.’’ Journal of Computing and Information Science in Engineering Vol. 23 No. 4 (2023): p. 041003.
  • [29] Ferrari, Alessio, Spagnolo, Giorgio Oronzo and Gnesi, Stefania. ‘‘Pure: A dataset of public requirements documents.’’ 2017 IEEE 25th International Requirements Engineering Conference (RE): pp. 502–505. 2017. IEEE.
  • [30] Chang, Yupeng, Wang, Xu, Wang, Jindong, Wu, Yuan, Zhu, Kaijie, Chen, Hao, Yang, Linyi, Yi, Xiaoyuan, Wang, Cunxiang, Wang, Yidong et al. ‘‘A survey on evaluation of large language models.’’ arXiv preprint arXiv:2307.03109 (2023).
  • [31] Papineni, Kishore, Roukos, Salim, Ward, Todd and Zhu, Wei-Jing. ‘‘Bleu: a method for automatic evaluation of machine translation.’’ Proceedings of the 40th annual meeting of the Association for Computational Linguistics: pp. 311–318. 2002.
  • [32] Blecher, Lukas, Cucurull, Guillem, Scialom, Thomas and Stojnic, Robert. ‘‘Nougat: Neural optical understanding for academic documents.’’ arXiv preprint arXiv:2308.13418 (2023).
  • [33] Lin, Chin-Yew. ‘‘Rouge: A package for automatic evaluation of summaries.’’ Text summarization branches out: pp. 74–81. 2004.
  • [34] Liu, Haotian, Li, Chunyuan, Wu, Qingyang and Lee, Yong Jae. ‘‘Visual Instruction Tuning.’’ (2023).
  • [35] Liu, Haotian, Li, Chunyuan, Li, Yuheng and Lee, Yong Jae. ‘‘Improved baselines with visual instruction tuning.’’ arXiv preprint arXiv:2310.03744 (2023).
  • [36] Liu, Jerry. ‘‘LlamaIndex.’’ (2022). 10.5281/zenodo.1234. URL https://github.com/jerryjliu/llama_index.

### v2

@@ 6 CONCLUSION
Using DesignQA, we conducted a rigorous evaluation on variants of several state-of-the-art (at the time of writing) MLLMs: GPT-4o, GPT-4, Claude-Opus, Gemini-1.0, and LLaVA-1.5. While GPT-4o was the best performing model of the MLLMs tested, we uncovered significant limitations in MLLMs’ current abilities to accurately interpret complex technical documents, specifically in referencing relevant requirements and in analyzing technical images. These findings illustrates that analyzing and designing according to technical documentation requires a diverse skill-set, and that there is a need to develop MLLMs better suited to the multifaceted-nature of engineering design. Our research highlights the need for advancements in AI models to enhance their comprehension of engineering requirements and documentation, suggesting directions for future efforts. Our work aims to bridge the gap in AI’s capability to support engineering design processes more effectively, paving the way for sophisticated AI-assisted engineering solutions.
@@ Acknowledgments
The authors wish to express their gratitude to Henry Smith and the MIT Motorsports team for their critical contribution to dataset creation and verification, which significantly enhanced the quality of this research. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2023345746. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
[…]
Appendix A: Model Parameters
All MLLMs evaluated by the authors were run through the LlamaIndex framework. We used the default parameters, set in LlamaIndex, for each model tested. For LLaVA-1.5, the default parameters are temperature of 0.75 and top_p of 0.9.666https://github.com/run-llama/llama_index/blob/main/llama-index-integrations/multi_modal_llms/llama-index-multi-modal-llms-replicate/llama_index/multi_modal_llms/replicate/base.py For GPT-4o, GPT-4, Gemini-1.0, and Claude-Opus, the default temperature set by the LlamaIndex framework is 0.1.777https://github.com/run-llama/llama_index/blob/main/llama-index-core/llama_index/core/constants.py For these models, LlamaIndex doesn’t set defaults for top_p, so each model should retain its own default top_p value. For the Rule Extraction questions, max_new_tokens was set to 250 for all models. For the Rule Comprehension questions, max_new_tokens was set to 100. For the Rule Compliance questions, max_new_tokens was set to 1500.
Appendix B: GuaranteedRAG
GuaranteedRAG for Retrieval Questions: For each Retrieval question, an 8716-character  [^8] Note that in generating the simple LlamaIndex RAG for most of the questions, except the Rule Compliance questions, the top 15 most relevant document chunks were used. These top 15 most relevant chunks had an average character count of 10895 characters. Initially, we tried generating 10895 character count excerpts for GuaranteedRAG, but these contexts, coupled with the prompts, often exceeded the token limit for LLaVA. This happened due to diffrences in character counts and token counts. So as to not run up against token limits, we reduced the GuaranteedRAG context length to 8716 characters for all questions. 8716 characters corresponds with the average length of the top 12 most relevant document chunks in the simple LlamaIndex RAG implementation. portion of the text was extracted from the full rulebook. This portion was specifically selected to contain the rule-in-question; the location of the rule-in-question within the excerpt was randomized.
GuaranteedRAG for Compilation Questions: The ground truth answer to each Compilation question is a list of rule numbers. For each Compilation question, we extracted the text for each of these ground truth rules from the rule document. If the total sum of these rule texts was less than 8716 characters, we randomly added in other rules until we reached the 8716 limit. If, on the other hand, the total sum of the ground truth rule texts was greater than 8716 before adding additional, random rules, we iteratively removed the longest rule from the list until we met the character limit. The rule texts were randomly shuffled (before appending them together) so that they were presented in no specific order.
GuaranteedRAG for Definition and Presence Questions: As explained in Section 3.1.3, each Definition and Presence question tests the model’s understanding of a vehicle component. The components are either: 1) mentioned explicitly in the definition section of the rule document (“definition component”), 2) mentioned multiple times throughout the document but are not explicitly in the definition section (“multi-mention component”), or 3) are not mentioned in the rule document at all (“no-mention component”). For each Definition or Presence question, we generate GuaranteedRAG in accordance with the component’s mention type. For definition component questions, we extract the two “Definition” sections of the rule document and randomly place them in between two other randomly selected segments of rule text, such that the total extracted text length is 8716 characters. For multi-mention component questions, we find every mention of the component within the rule document. For each mention, we extract a constant length text excerpt from the rule document with the mention centered in the excerpt. When appended together, the total length of these text excerpts is 8716 characters. For no-mention component questions, we randomly select an 8716-character segment from the rule document.
GuaranteedRAG for Dimension and Functional Performance Questions: Each of these questions asks the model whether a design conforms with a specific rule number. For each of these questions, an 8716-character portion of the text was extracted from the rulebook. This portion is guaranteed to contain the rule text of the rule in question; the location of the rule-in-question within the excerpt was randomized. In some cases, the rule in question references another rule in the FSAE rule document. In these cases, the GuaranteedRAG is split into two even chunks that are appended together to form the full 8716-character GuaranteedRAG: the first half is guaranteed to contain the rule text for the rule (at a random location) and the second half is guaranteed to contain the rule text for the referenced rule (at a random location).